# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v2

@@ 4 Case Studies
We evaluate and benchmark UA-SABI against its uncertainty-unaware counterpart SABI, the corresponding MCMC-based methods, and (if possible) standard ABI using a toy example and two real-world case studies. All code and data will be made publicly available upon acceptance.
@@ 4.1 Objectives
We want to show the advantage of using surrogate data instead of scarce simulation data, and demonstrate the importance of quantifying and propagating the uncertainty of the surrogate model. To do so, we compare UA-SABI, SABI, and standard ABI for the true model – trained with full and low budget (shown in blue in Fig.˜2). To validate the posteriors produced by these methods, we also compare them with those obtained using the corresponding MCMC-based approach (shown in purple in Fig.˜2). Further, to highlight the efficiency of our method, we compare the runtimes of UA-SABI and E-Post (shown in red in Fig.˜2).
To assess posterior calibration for our first two case studies, we generate multiple synthetic ground truth parameters and perform inference on the simulated datasets. For each dataset, we compute the rank of the ground truth within the posterior samples and summarize the results using empirical cumulative distribution function (ECDF) difference plots – a procedure known as simulation-based calibration (SBC) checking (Talts et al., 2020; Säilynoja et al., 2022; Modrák et al., 2023). To compare runtimes, we estimated how many observation sets are needed before the accumulated computational cost of repeated E-Post inference runs exceeds the total cost of UA-SABI, including both training and its repeated (quasi-instant) inference.
@@
For all case studies, we use polynomial chaos expansion (PCE) for the surrogate model (Wiener, 1938; Sudret, 2008; Oladyshkin & Nowak, 2012), as it is a fast-to-evaluate and fast-to-train surrogate and a flexible black-box model which can be applied to (almost) arbitrary simulation models. In principle, PCE is just one example among many possible choices – any uncertainty-aware surrogate could be used and replaced in a modular fashion depending on the application domain. A deterministic PCE constructs the surrogate through a spectral projection onto orthogonal (w.r.t. $p(\mathbf{x},\bm{\omega})$) polynomial basis functions, expressed as
@@
To construct a Bayesian PCE (Shao et al., 2017; Bürkner et al., 2023), we define a prior distribution over the surrogate coefficients $p(\mathbf{c})$ and the approximation error parameter $p(\sigma$). We choose a normal likelihood:
$p(\mathbf{y}\mid\mathbf{x},\bm{\omega},\mathbf{c},\sigma)=\mathcal{N}(M(\mathbf{x},\bm{\omega})\mid\widetilde{M}_{\mathbf{c}}(\mathbf{x},\bm{\omega}),\sigma).$ (12)
Employing Hamiltonian Monte Carlo (Hoffman & Gelman, 2014) yields samples from the surrogate posterior $p(\mathbf{c},\sigma\mid D_{T})$ (Bürkner et al., 2023). The resulting Bayesian PCEs serve as our uncertainty-aware surrogate.
@@ 4.2 Case Study 1: LogSin Model
$y=M(x,\omega)=\omega\log(x)+\sin(0.05x)+0.01x+1.$ (13)
@@ 4.2.1 Setup
To perform inference on observation sets, we sample 200 ground truth parameters $\omega^{*}$ from a prior $p(\omega)=\mathcal{N}(1,0.2)$ and generate four $(\mathbf{x},\mathbf{y})$ observations for each. We compare the inference results of our proposed surrogate-based ABI to surrogate-based MCMC methods. For ABI, we used an equivariant deep set summary network (Zaheer et al., 2017) and a coupling flow inference network. We employ online training, where newly sampled surrogate outputs are used at each iteration. All ABI models were trained for 100 epochs, with a batch size of 64 and 128 batches per epoch, using BayesFlow (Radev et al., 2023b). For given observations, ABI generates 4,000 posterior samples via the inference network. Surrogate-based MCMC (Point) draws 4,000 samples using 4 chains (1,000 warm-up and 1,000 sampling iterations). For E-Post, we run MCMC separately for each surrogate posterior draw (1,000 warm-up, 4 sampling iterations per run), a process that is embarrassingly parallelizable but still computationally expensive. Further computational details are given in Appendix˜C.
@@ 4.2.2 Impact of Uncertainty Propagation
First, we compare the performance of our two surrogate-based methods, SABI and UA-SABI, against each other. Then we compare them against a low-budget ABI, trained with the same simulation data as used for surrogate training, and a full-budget ABI trained on sufficient simulation data. Figure˜3(a) shows corresponding recovery plots. They show the posterior median (circles) and median deviation (vertical lines) for any given ground truth value $\omega^{*}$.
@@
Figure˜3(a) shows that a low-budget ABI model is unable to recover the ground truth values, particularly towards the boundaries of the parameter domain. Using a surrogate as part of SABI or UA-SABI, it becomes possible to generate sufficient ABI training data. Assigning the inter- and extrapolation tasks to the surrogate, specifically designed for this purpose, produces better outcomes than relying on the NPE to implicitly interpolate the posterior from scarce data. However, without propagating surrogate uncertainty (i.e., using SABI), estimated posterior uncertainty remains minimal and does not cover the true values, thus producing overconfident posteriors. Conversely, incorporating surrogate uncertainty via UA-SABI leads to an increase in posterior uncertainty, now better covering the true values. Indeed, the ECDF difference plots showing calibration in Fig.˜3(b) confirm that both low-budget ABI and SABI produce severely miscalibrated posteriors, while UA-SABI and full-budget ABI yield well-calibrated posteriors. However, achieving this calibration with full-budget ABI requires substantially more training data, highlighting the efficiency of UA-SABI under limited simulation budgets.
@@ 4.2.3 Validation of Parameter Posterior
We validate the correctness of the surrogate-based ABI approaches and benchmark their performance against a reference solution. To this end, we compare the two surrogate-based ABI methods, SABI and UA-SABI, against the corresponding surrogate-based MCMC methods, Point and E-Post. We plot the posterior medians (circles) and their median deviations (lines) for the ABI methods along with the corresponding MCMC results, as well as the ECDF differences for all the methods in Fig.˜4.
@@
When comparing the ABI posteriors to the MCMC full reference solution, we observe that for all methods, the ABI results align closely with their MCMC counterparts. This validates the reliability and confirms the correctness of the ABI methods. Comparing the ECDF differences between the ABI methods and their corresponding MCMC baselines shows that both Point and SABI produce similarly overconfident posteriors, while UA-SABI yields well-calibrated posteriors. In contrast, E-Post tends to produce even slightly underconfident posteriors.
@@ 4.2.4 Runtime Comparison
Next, we justify the computational effort required to train UA-SABI. Specifically, we aim to determine the break-even point – that is, the number of observation sets (i.e., measurement series) after which training an UA-SABI and performing (quasi-instant) inference becomes more efficient than repeatedly rerunning E-Post. Beyond that point, the one-time training cost of UA-SABI is amortized. All experiments were performed on a standard laptop equipped with an Intel Core i7-1185G7 CPU. For ABI, we consider both the upfront training phase and the inference time for the given sets as part of the total runtime. For E-Post, we measure the inference runtime while parallelized across 8 cores. Unlike for UA-SABI, E-Post runtimes depend on the number of cores used for parallelization.
In Fig.˜6(a), we show the runtime of UA-SABI and E-Post for $\{5,10,15,20\}$ inference runs. We observe that UA-SABI’s runtime is nearly constant for the number of inference runs, while E-Post scales linearly. Based on this comparison, UA-SABI is already justified after around 9 inference runs.
@@ 4.3.1 Setup
In general, the setup for the $\text{CO}_{2}$ storage model follows the same structure as that described in Section˜4.2.1. The priors of the input parameters are given as $\text{IR}\sim 6.4\times 10^{-4}\times(1+\text{Beta}(4,2))$, $\text{PM}\sim 2\times\text{Beta}(1.25,1.25)+2$, and $\text{PR}\sim\text{Beta}(2.4,9)$ (Köppel et al., 2019). For PCE training, we considered 64 Sobol sequence evaluations scaled to the input parameter prior ranges and constructed the polynomial basis with arbitrary polynomial chaos (aPC) based on the priors (Oladyshkin & Nowak, 2012; Bürkner et al., 2023).
@@ 4.3.2 Validation of Parameter Posterior
We compare the inferred posterior samples obtained from a full-budget ABI ($N_{B}=10^{4}$), a low-budget ABI ($N_{B}=64$), SABI, and UA-SABI. Full-budget ABI was feasible due to the already available data (Köppel et al., 2017; Oladyshkin et al., 2020). Therefore, recovery plots were chosen to validate the quality of our results relative to standard ABI (Fig.˜5).
Recovery plots in Fig.˜5(a) again show that low-budget ABI struggles to recover the ground truth parameter, particularly near domain boundaries. SABI also produces poor estimates and, moreover, fails to capture uncertainty in the inferred parameter posteriors. This overconfidence results in miscalibrated posteriors, confirmed by the two corresponding ECDF difference plots in Fig.˜5(b).
In contrast, UA-SABI performs comparably to full-budget ABI while accounting for the additional uncertainty introduced by using a surrogate to generate training data. According to the ECDF difference plots, it produces a well-calibrated posterior for porosity. Additionally, recovery plots between ABI and MCMC for surrogate-based methods and ECDF difference plots for MCMC-based methods are provided in Appendix˜D. Despite the model’s discontinuities and the resulting surrogate misspecification, our method successfully quantifies and propagates surrogate uncertainty to the parameter posterior.
@@ 4.3.3 Runtime Comparison
Figure˜6(b) presents the measured runtimes of UA-SABI and E-Post for $\{5,6,7,8\}$ inference runs, whereby E-Post was parallelized on 16 cores. We observe a break-even point between 6 and 7 inference runs, indicating that UA-SABI becomes justified after 7 runs. Despite using more cores for E-Post, the training costs are amortized earlier for a more expensive model.
@@ 4.4 Case Study 3: Microbially Induced Calcite Precipitation Model
In this case study, we transition from benchmarks to realistic conditions by using real measurement data. Case studies 1 and 2 introduced a toy example and a virtual benchmark, respectively. Case study 3 now focuses on posterior inference from observed measurements, where the primary goal is to evaluate the applicability of UA-SABI under real-world conditions. A runtime comparison for the MICP model on synthetic data is added in Section˜E.4.
The underlying model predicts microbially induced calcite precipitation (MICP). MICP is a reactive transport process, here taking place inside a porous medium. It includes at least biofilm, calcite, and the unreactive solid matrix as solid phases, water, and, in some cases, an additional fluid phase such as gas. Relevant in the aqueous phase are dissolved calcium, inorganic carbon, and urea. The complete set of relevant species is detailed in Hommel et al. (2015). The process is mediated by S. pasteurii, a bacterium that produces the enzyme urease. Urease catalyzes the hydrolysis of urea into ammonia and carbonic acid. In aqueous solution, ammonia consumes hydrogen ions ($H^{+}$), which raises the pH. The increase in pH alters carbonate equilibria: carbonic acid dissociates further, releasing additional $H^{+}$ and carbonate ions. This raises the concentration of dissolved carbonate species. When calcium ions are present, they combine with carbonate ions and trigger the precipitation of calcite within the pore space.
The experiment is performed on a $61\text{cm}$ sand-filled column with a $2.54\text{cm}$ diameter. Bacteria are first injected from the bottom, followed by an overnight no-flow period to allow biofilm formation. Biofilm growth is then stimulated with a $24\text{h}$ substrate injection. Afterwards, two pore volumes of a $0.33\text{mol}/\text{L}$ calcium–urea solution are injected at $10\text{ml}/\text{min}$, repeated every $24\text{h}$. Each injection is followed by a no-flow period for mineralization, and then a substrate injection to reactivate the biofilm (Hommel et al., 2015). This cycle is repeated 30 times. A more detailed description can be found in Hommel et al. (2015).
@@ 4.4.1 Setup
In general, the setup for the MICP model follows the same structure as that described in the former case studies. The priors of the input parameters are given as $c_{a,1}\sim\mathcal{U}(10^{-10},10^{-7})$, $c_{a,2}\sim\mathcal{U}(10^{-10},10^{-6})$, $\rho_{f}\sim\mathcal{U}(1,15)$, and $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-4})$ (Scheurer et al., 2021). For PCE training, 25 model evaluations within the prior ranges of input parameters are available, which are not necessarily optimal for surrogate training. We constructed the polynomial basis with aPC based on the available model evaluations (Oladyshkin & Nowak, 2012; Bürkner et al., 2023).

### v3

@@ 4 Case Studies
We evaluate and benchmark UA-SABI against its uncertainty-unaware counterpart SABI, the corresponding MCMC-based methods, and (if possible) standard ABI using a toy example and two real-world case studies. Code is publicly available on GitHub [^1] https://github.com/LS3-university-of-stuttgart/ua-sabi-paper.
@@ 4.1 Objectives
We want to show the advantage of using surrogate data instead of scarce simulation data, and demonstrate the importance of quantifying and propagating the uncertainty of the surrogate model. To do so, we compare UA-SABI, SABI, and standard ABI for the true model – trained with full and low budget (shown in blue in Fig.˜3). To validate the posteriors produced by these methods, we also compare them with those obtained using the corresponding MCMC-based approach (shown in purple in Fig.˜3). Further, to highlight the efficiency of our method, we compare the runtimes of UA-SABI and E-Post (shown in red in Fig.˜3).
To assess posterior calibration for our first two case studies, we generate multiple synthetic ground truth parameters and perform inference on the simulated datasets. For each dataset, we compute the rank of the ground truth within the posterior samples and summarize the results using empirical cumulative distribution function (ECDF) difference plots – a procedure known as simulation-based calibration (SBC) checking (Talts et al., 2020; Säilynoja et al., 2022; Modrák et al., 2023). SBC assesses whether the inference method is calibrated by verifying that the rank of the true parameter among posterior samples follows a uniform distribution across repeated simulations. To compare runtimes, we estimated how many observation sets are needed before the accumulated computational cost of repeated E-Post inference runs exceeds the total cost of UA-SABI, including both training and its repeated (quasi-instant) inference.
@@
For all case studies, we use polynomial chaos expansion (PCE) for the surrogate model (Wiener, 1938; Sudret, 2008; Oladyshkin and Nowak, 2012), as it is a fast-to-evaluate and fast-to-train surrogate and a flexible black-box model which can be applied to (almost) arbitrary simulation models. In principle, PCE is just one example among many possible choices – any uncertainty-aware surrogate could be used and replaced in a modular fashion depending on the application domain. A deterministic PCE constructs the surrogate through a spectral projection onto orthogonal (w.r.t. $p(\mathbf{x},\bm{\omega})$) polynomial basis functions, expressed as
@@
To construct a Bayesian PCE (Shao et al., 2017; Bürkner et al., 2023), we define a prior distribution over the surrogate coefficients $p(\mathbf{c})$ and the approximation error parameter $p(\sigma$). We choose a normal likelihood for the simulator output $\mathbf{y}=M(\mathbf{x},\bm{\omega})$:
$p(\mathbf{y}\mid\mathbf{x},\bm{\omega},\mathbf{c},\sigma)=\mathcal{N}(\mathbf{y}\mid\widetilde{M}_{\mathbf{c}}(\mathbf{x},\bm{\omega}),\sigma).$ (15)
Employing Hamiltonian Monte Carlo (Hoffman and Gelman, 2014) yields samples from the surrogate posterior $p(\mathbf{c},\sigma\mid D_{T})$ (Bürkner et al., 2023). The resulting Bayesian PCEs serve as our uncertainty-aware surrogate.
@@ 4.2 Case Study 1: LogSin Model
$y=M(\text{x},\omega)=\omega\log(\text{x})+\sin(0.05\text{x})+0.01\text{x}+1.$ (16)
@@ 4.2.1 Setup
To perform inference on observation sets, we sample 200 ground truth parameters $\omega^{*}$ from a prior $p(\omega)=\mathcal{N}(1,0.2)$ and generate four $(\mathbf{x},\mathbf{y})$ observations for each. We compare the inference results of our proposed surrogate-based ABI to surrogate-based MCMC methods. For ABI, we used an equivariant DeepSet summary network (Zaheer et al., 2017), given $p(\mathbf{y}\mid\mathbf{x})$ being i.i.d., and a coupling flow inference network. We employ online training, where newly sampled surrogate outputs are used at each iteration. All ABI models were trained for 100 epochs, with a batch size of 64 and 128 batches per epoch, using BayesFlow (Radev et al., 2023b). For given observations, ABI generates 4,000 posterior samples via the inference network. Surrogate-based MCMC (Point) draws 4,000 samples using 4 chains (1,000 warm-up and 1,000 sampling iterations). For E-Post, we run MCMC separately for each surrogate posterior draw (1,000 warm-up, 4 sampling iterations per run), a process that is embarrassingly parallelizable but still computationally expensive. Further computational details are given in Appendix˜C.
@@ 4.2.2 Impact of Uncertainty Propagation
First, we compare the performance of our two surrogate-based methods, SABI and UA-SABI, against each other. Then we compare them against a low-budget ABI, trained with the same simulation data as used for surrogate training, and a full-budget ABI trained on sufficient simulation data. Figure˜4(a) shows corresponding recovery plots. They show the posterior median (circles) and median deviation (vertical lines) for any given ground truth value $\omega^{*}$.
@@
Figure˜4(a) shows that a low-budget ABI model is unable to recover the ground truth values, particularly towards the boundaries of the parameter domain. Using a surrogate as part of SABI or UA-SABI, it becomes possible to generate sufficient ABI training data. Assigning the inter- and extrapolation tasks to the surrogate, specifically designed for this purpose, produces better outcomes than relying on the NPE to implicitly interpolate the posterior from scarce data. However, without propagating surrogate uncertainty (i.e., using SABI), estimated posterior uncertainty remains minimal and does not cover the true values, thus producing overconfident posteriors. Conversely, incorporating surrogate uncertainty via UA-SABI leads to an increase in posterior uncertainty, now better covering the true values. Indeed, the ECDF difference plots showing calibration in Fig.˜4(b) confirm that both low-budget ABI and SABI produce severely miscalibrated posteriors, while UA-SABI and full-budget ABI yield well-calibrated posteriors. However, achieving this calibration with full-budget ABI requires substantially more training data, highlighting the efficiency of UA-SABI under limited simulation budgets.
@@ 4.2.3 Validation of Parameter Posterior
We validate the correctness of the surrogate-based ABI approaches and benchmark their performance against a reference solution. To this end, we compare the two surrogate-based ABI methods, SABI and UA-SABI, against the corresponding surrogate-based MCMC methods, Point and E-Post. We plot the posterior medians (circles) and their median deviations (lines) for the ABI methods along with the corresponding MCMC results, as well as the ECDF differences for all the methods in Fig.˜5.
@@
When comparing the ABI posteriors to the MCMC full reference solution, we observe that for the uncertainty-unaware methods, the SABI results align closely with the MCMC counterpart. Comparing the respective ECDF differences it shows that both Point and SABI produce similarly overconfident posteriors. For the uncertainty-aware methods, E-Post and UA-SABI, we observe a slight difference in the recovery plots and the ECDF differences. To investigate this discrepancy, we performed a convergence analysis for the MCMC sampling of E-Post, with results shown in Appendix˜D. The analysis confirms that the difference can be attributed to non-convergence of many MCMC runs. Examining the calibration results reveals that while E-Post is slightly underconfident, UA-SABI yields well-calibrated posteriors. This indicates the reliability and correctness of UA-SABI.
@@ 4.2.4 Runtime Comparison
Next, we justify the computational effort required to train UA-SABI. Specifically, we aim to determine the break-even point – that is, the number of observation sets (i.e., measurement series) after which training UA-SABI and performing (quasi-instant) inference becomes more efficient than repeatedly rerunning E-Post. Beyond that point, the one-time training cost of UA-SABI is amortized. All experiments were performed on a standard laptop equipped with an Intel Core i7-1185G7 CPU. For ABI, we consider both the upfront training phase and the inference time for the given sets as part of the total runtime. For E-Post, we measure the inference runtime while parallelized across 8 cores. Unlike for UA-SABI, E-Post runtimes depend on the number of cores used for parallelization.
In Fig.˜7(a), we show the runtime of UA-SABI and E-Post for $\{5,10,15,20\}$ inference runs. We observe that UA-SABI’s runtime is nearly constant for the number of inference runs, while E-Post scales linearly. Based on this comparison, UA-SABI is already justified after around 9 inference runs.
@@ 4.3.1 Setup
In general, the setup for the $\text{CO}_{2}$ storage model follows the same structure as that described in Section˜4.2.1. The priors of the input parameters are given as $\text{IR}\sim 6.4\times 10^{-4}\times(1+\text{Beta}(4,2))$, $\text{PM}\sim 2\times\text{Beta}(1.25,1.25)+2$, and $\text{PR}\sim\text{Beta}(2.4,9)$ (Köppel et al., 2019). For PCE training, we considered 64 Sobol sequence evaluations scaled to the input parameter prior ranges and constructed the polynomial basis with arbitrary polynomial chaos (aPC) based on the priors (Oladyshkin and Nowak, 2012; Bürkner et al., 2023).
@@ 4.3.2 Validation of Parameter Posterior
We compare the inferred posterior samples obtained from a full-budget ABI ($N_{B}=10^{4}$), a low-budget ABI ($N_{B}=64$), SABI, and UA-SABI. Full-budget ABI was feasible due to the already available data (Köppel et al., 2017; Oladyshkin et al., 2020). Therefore, recovery plots were chosen to validate the quality of our results relative to standard ABI (Fig.˜6).
Recovery plots in Fig.˜6(a) again show that low-budget ABI struggles to recover the ground truth parameter, particularly near domain boundaries. SABI also produces poor estimates and, moreover, fails to capture uncertainty in the inferred parameter posteriors. This overconfidence results in miscalibrated posteriors, confirmed by the two corresponding ECDF difference plots in Fig.˜6(b).
In contrast, UA-SABI performs comparably to full-budget ABI while accounting for the additional uncertainty introduced by using a surrogate to generate training data. According to the ECDF difference plots, it produces a well-calibrated posterior for porosity. Additionally, recovery plots between ABI and MCMC for surrogate-based methods and ECDF difference plots for MCMC-based methods are provided in Appendix˜E. Despite the model’s discontinuities and the resulting surrogate misspecification, our method successfully quantifies and propagates surrogate uncertainty to the parameter posterior.
@@ 4.3.3 Runtime Comparison
Figure˜7(b) presents the measured runtimes of UA-SABI and E-Post for $\{5,6,7,8\}$ inference runs, whereby E-Post was parallelized on 16 cores. We observe a break-even point between 6 and 7 inference runs, indicating that UA-SABI becomes justified after 7 runs. Despite using more cores for E-Post, the training costs are amortized earlier for a more expensive model.
@@ 4.4 Case Study 3: Microbially Induced Calcite Precipitation Model
In this case study, we transition from benchmarks to realistic conditions by using real measurement data. Case studies 1 and 2 introduced a toy example and a virtual benchmark, respectively. Case study 3 now focuses on posterior inference from observed measurements, where the primary goal is to evaluate the applicability of UA-SABI under real-world conditions. Specifically, in addition to simulation-to-surrogate misspecification, we encounter simulation-to-real and thus surrogate-to-real misspecification. A runtime comparison for the MICP model on synthetic data is added in Section˜F.5.
The underlying model predicts microbially induced calcite precipitation (MICP). MICP is a reactive transport process, here taking place inside a porous medium. It includes at least biofilm, calcite, and the unreactive solid matrix as solid phases, water, and, in some cases, an additional fluid phase such as gas. Relevant in the aqueous phase are dissolved calcium, inorganic carbon, and urea. The complete set of relevant species is detailed in Hommel et al. (2015). The process is mediated by S. pasteurii, a bacterium that produces the enzyme urease. Urease catalyzes the hydrolysis of urea into ammonia and carbonic acid. In aqueous solution, ammonia consumes hydrogen ions ($H^{+}$), which raises the pH. The increase in pH alters carbonate equilibria: carbonic acid dissociates further, releasing additional $H^{+}$ and carbonate ions. This raises the concentration of dissolved carbonate species. When calcium ions are present, they combine with carbonate ions and trigger the precipitation of calcite within the pore space. A schematic illustration of the relevant process during MICP is given in Section˜F.1.
The experiment is performed on a $61\text{cm}$ sand-filled column with a $2.54\text{cm}$ diameter. Bacteria are first injected from the bottom, followed by an overnight no-flow period to allow biofilm formation. Biofilm growth is then stimulated with a $24\text{h}$ substrate injection. Afterwards, two pore volumes of a $0.33\text{mol}/\text{L}$ calcium-urea solution are injected at $10\text{ml}/\text{min}$, repeated every $24\text{h}$. Each injection is followed by a no-flow period for mineralization, and then a substrate injection to reactivate the biofilm (Hommel et al., 2015). This cycle is repeated 30 times. A more detailed description can be found in Hommel et al. (2015).
@@ 4.4.1 Setup
In general, the setup for the MICP model follows the same structure as that described in the former case studies. The priors of the input parameters are given as $c_{a,1}\sim\mathcal{U}(10^{-10},10^{-7})$, $c_{a,2}\sim\mathcal{U}(10^{-10},10^{-6})$, $\rho_{f}\sim\mathcal{U}(1,15)$, and $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-4})$ (Scheurer et al., 2021). For PCE training, 25 model evaluations within the prior ranges of input parameters are available, which are not necessarily optimal for surrogate training. We constructed the polynomial basis with aPC based on the available model evaluations (Oladyshkin and Nowak, 2012; Bürkner et al., 2023).